OpenStudy (thats_kyy):

Jolie can read around 14 sentences per minute of her current novel. She wants to finish her reading before the movie comes out next week. If she has 17 pages left to read and each page has about 120 sentences, how many hours should Jolie set aside to complete her reading?

OpenStudy (thats_kyy):

@mathstudent55 @Nnesha

OpenStudy (thats_kyy):

Nooo no this

OpenStudy (thats_kyy):

Jolie can read around 14 sentences per minute of her current novel. She wants to finish her reading before the movie comes out next week. If she has 17 pages left to read and each page has about 120 sentences, how many hours should Jolie set aside to complete her reading? Show your work to receive full credit. (10 points)

OpenStudy (thats_kyy):

i mean this one

Nnesha (nnesha):

substitute g(x) for 20 then solve for x \[\huge\rm 20=2(x-4)\]

OpenStudy (thats_kyy):

@Nnesha

OpenStudy (thats_kyy):

i dont want an answer i just need understanding

OpenStudy (mathstudent55):

Let's calculate it minutes first, then we will convert the answer to hours.

OpenStudy (thats_kyy):

okay

OpenStudy (thats_kyy):

Soo Hours=60minutes

OpenStudy (mathstudent55):

If she has 17 pages left, and each page has 120 sentences, how many sentences total are there left to read?

OpenStudy (mathstudent55):

Yes. We will use 1 hour = 60 minutes at the end to convert to hours, but let's find the number of minutes first.

OpenStudy (thats_kyy):

2040

OpenStudy (thats_kyy):

2040 sentences

OpenStudy (mathstudent55):

Good. She has to read 2040 sentences. She takes 1 minute to read 14 sentences. 2040 sentences is how many times 14 sentences?

OpenStudy (thats_kyy):

28560

OpenStudy (thats_kyy):

no

OpenStudy (thats_kyy):

wait

OpenStudy (mathstudent55):

No. The other way around. This time you need to divide.

OpenStudy (thats_kyy):

okay

OpenStudy (thats_kyy):

145.7

OpenStudy (thats_kyy):

@mathstudent55

OpenStudy (mathstudent55):

Good. Since 2040 is 145.7 times 14, that means she needs 145.7 times 1 minute. That means she needs 145.7 minutes.

OpenStudy (mathstudent55):

Now we convert minutes to hours.

OpenStudy (thats_kyy):

so divide 145.7/60

OpenStudy (thats_kyy):

@mathstudent55

OpenStudy (thats_kyy):

@Nnesha

OpenStudy (thats_kyy):

may you please help me out

OpenStudy (thats_kyy):

what do you mean?

OpenStudy (mathstudent55):

Correct. Divide 145.7 by 60 to get hours.

OpenStudy (thats_kyy):

okay so i do i write and explanation

OpenStudy (thats_kyy):

do i write the steps

OpenStudy (mathstudent55):

You write: She needs to read 17 pages with 120 sentences in each page. 17 * 120 = 2040 She needs to read 2040 sentences. She reads 14 sentences in 1 minute. 2040/14 = 145.7 minutes. She needs 145.7 minutes. 1 hour = 60 minutes 145.7/60 = 2.4 hours

OpenStudy (thats_kyy):

Yoooooo Thamksss

OpenStudy (thats_kyy):

Now I understand It

OpenStudy (mathstudent55):

Great. You're welcome.